OpenStudy (anonymous):

limit of n^x/x! as x goes to inf is 0 ,can u explain y?

OpenStudy (anonymous):

Looks like a taylor series expansion.

OpenStudy (anonymous):

Did you know \[ e^x = \sum_{n=0}^{\infty }\frac{x^n}{n!} \]

OpenStudy (anonymous):

i know the limit of 1/x! will be zero but what about n^x as x-->inf?

OpenStudy (anonymous):

Do you know l'hospital's rule?

OpenStudy (anonymous):

What you need to realize is that the \(x!\) and the \(n^x\) are both competing.

OpenStudy (anonymous):

yes,u mean Derivation of the numerator and denominator

OpenStudy (anonymous):

what the limit of n^x? is that correct ,in this question we have inf/inf so ,we use l'hospital's rule?

OpenStudy (anonymous):

We have an indeterminate form \(\infty / \infty\)